gwaylat:

Sonji paid $25 for two scarves, which were different prices. When she got home she could not find the receipt. She remembered that one scarf cost $3 more than the other. What was the price of the more expensive scarf?

crispyrat:

let x be the price of one scarf that cost 3 more and y the other one. If x is 3 more than that means y+3 is the same as x so y+3=x Now we know that x+y=25 so we plug in the equation and get 2y+3=25 now just solve the equation and plug the y value into y+3 and you will get x along with y